OpenStudy (samara89):

What is the base 5 representation of 30?

OpenStudy (wolf1728):

To convert from base 10 to another base (Base 5), you divide 30 / 5 = 6 remainder of 0 (rightmost number) 6 / 5 = 1 remainder of 1 (middle number) 1 / 5 = 0 remainder of 1 (leftmost number) So 30 in base 5 = 110 To check the result you multiply the rightmost digit by 1 multiply the next digit by 5 multiply the next digit by 25 Then add the results 0 * 1 = 0 1 * 5 = 5 1 * 25 = 25 This totals 30